5 . Which of the following equations contains the solution points given?
x 1 5 10 9
y 450 90 45 50
A. xy = k
B. x2 + k = y
C. 4x – k = y
D. y = kx – 5

Answer by checkley75(3666)

xy=k
1*450=450
5*90=450
10*45=450
9*50=450
k being a constant=450